2 questions:
1. If MO=8, PO=12, AND MP=4, which point lies between the other two?
~ My answer was point M
2. If angle EFG is a right angle
angle Efh= x+y
angle HFG= 2x-6
solve for y in terms of x
~my answer was y= x-6
please reply asap!
_____________________

Answer by jainenderkapoor(61)

(a) you are correct M lies in between P and O
MO = 8 PO = 12 and MP = 4
12 = 8 + 4
we can write PO = PM + MO which shows that M is in between P and O
(b)
H lies in between EFG and F is the common vertex
so we get EFH + HFG = EFG
x + y + 2x - 6 = 90
3x + y - 6 = 90
y = 90 - 3x + 6
y = 96 - 3x

I hope this is clear to you.
If you need more clarifications, you can contact me at geniusy_2006@yahoo.com
I am also available for Online Tutoring at reasonable rates.